OpenStudy (cheesecakekitten):

WILL MEDAL AND FAN!!!!! What is the slope of a line perpendicular to the line with equation y = 5x - 2? A. 5 B. -1/5 C. -2 D. -5

OpenStudy (smilie2015):

B bcz for slope we use y=mx+c here m1 is 5 and for perpendicular m1=-1/m2 5=-1/5